OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (aroub):

y=mx+b , m=7 , (1,-3) -3=7(1)+b -3=7+b b=-10 => y=7x-10